OpenStudy (anonymous):

Using elimination to solve the system 2x + y = 8 and x + y = 4, which of the following could be the first step? A. 2x + y = 8--------> 2x + y = 8 (x + y = 4)(-1)---> -x - y = -4 B. 2x + (4 - x) = 8 C. 2x + y = x + y D. 3x + 2y = 12 E. x = 4

OpenStudy (vishal_kothari):

x—2 equations combined: 2x + y - x - y = 8 - 4 x = 4 y—1st equation, substitute x with 4: 2(4) + y = 8 8 + y = 8 y = 0 Answer: x = 4, y = 0 Proof—2nd equation: 4 + 0 = 4 4 = 4....

OpenStudy (anonymous):

So the answer is ?

OpenStudy (vishal_kothari):

Answer: x = 4, y = 0

OpenStudy (vishal_kothari):

i mean x = 4

OpenStudy (anonymous):

but its asking for the first step

OpenStudy (vishal_kothari):

i also that it is the first step....

OpenStudy (vishal_kothari):

*mean

OpenStudy (anonymous):

ok

OpenStudy (vishal_kothari):

thanks...